# Infinite-Dimensional Stochastic Differential Equations and Diffusion Dynamics of Coulomb Random Point Fields

### v3

@@ 8. Weak solutions of ISDE for the lower Dirichlet form
Recall that $\underline{\mathbf{X}}_{R}$ is a weak solution of SDEs (2.58) and (2.59). Hence, from Corollary 2, the diffusion $\underline{\mathbf{X}}_{R,\mathsf{y}}^{[m]}$, $\mathsf{y}\in\mathsf{S}$, given by $(\underline{\mathscr{E}}_{R\star}^{\mu^{[m]}},\underline{\mathscr{D}}_{R\star}^{\mu^{[m]}})$ and $\underline{\mathbf{X}}_{R}^{[m]}$ given by $(\mathscr{E}^{\mu^{[m]}}_{R},\underline{\mathscr{D}}_{R}^{\mu^{[m]}})$ coincide. This implies the identity of the Dirichlet forms in (11.2). ∎
Let $\mathscr{D}_{\bullet}^{\mu^{[m]}}=\{f\in\mathscr{D}_{\bullet}^{[m]};\mathscr{E}^{\mu^{[m]}}(f,f)<\infty,f\in L^{2}(\mu^{[m]})\}$. Let $(\mathscr{E}^{\mu^{[m]}},\underline{\mathscr{D}}^{\mu^{[m]}}_{\bullet})$ be the closure of $(\mathscr{E}^{\mu^{[m]}},\mathscr{D}_{\bullet}^{\mu^{[m]}})$ on $L^{2}(\mu^{[m]})$ given by Lemma 7.4.
@@
From (7.11), $(\mathscr{E}^{\mu^{[m]}}_{R},\underline{\mathscr{D}}_{R}^{\mu^{[m]}})$ is increasing. Hence, $(\underline{\mathscr{E}}_{R\star}^{\mu^{[m]}},\underline{\mathscr{D}}_{R\star}^{\mu^{[m]}})$ is increasing from Lemma 11.1. Let $(\underline{\mathscr{E}}_{\star}^{\mu^{[m]}},\underline{\mathscr{D}}_{\star}^{\mu^{[m]}})$ be the increasing limit of $(\underline{\mathscr{E}}_{R\star}^{\mu^{[m]}},\underline{\mathscr{D}}_{R\star}^{\mu^{[m]}})$.
@@
Assume ((A1))((A3)). Then, for each $m\in\{0\}\cup\mathbb{N}$,
$(\mathscr{E}^{\mu^{[m]}},\underline{\mathscr{D}}^{\mu^{[m]}}_{\bullet})$ $=(\mathscr{E}^{\mu^{[m]}},\underline{\mathscr{D}}^{\mu^{[m]}})=(\underline{\mathscr{E}}_{\star}^{\mu^{[m]}},\underline{\mathscr{D}}_{\star}^{\mu^{[m]}}).$ (11.3)
@@
The first equality in (11.3) follows from Lemma 7.5 i. The second follows from
$(\mathscr{E}^{\mu^{[m]}},\underline{\mathscr{D}}^{\mu^{[m]}})=\lim_{R\to\infty}(\mathscr{E}^{\mu^{[m]}}_{R},\underline{\mathscr{D}}_{R}^{\mu^{[m]}})=\lim_{R\to\infty}(\underline{\mathscr{E}}_{R\star}^{\mu^{[m]}},\underline{\mathscr{D}}_{R\star}^{\mu^{[m]}})=(\underline{\mathscr{E}}_{\star}^{\mu^{[m]}},\underline{\mathscr{D}}_{\star}^{\mu^{[m]}}).$
Here, we used Lemma 7.3 i and Lemma 11.1. ∎
@@
It is plausible that the following identity holds:
$(\mathscr{E}^{\mu^{[m]}},\overline{\mathscr{D}}^{\mu^{[m]}})=(\mathscr{E}^{\mu^{[m]}},\underline{\mathscr{D}}^{\mu^{[m]}}).$ (11.4)
Kawamoto proved (11.4) for Dyson’s model [14].
Let $(\mathscr{E}_{R}^{\mu_{\mathsf{a},R,\mathsf{y}}^{[m]}},\underline{\mathscr{D}}_{R}^{\mu_{\mathsf{a},R,\mathsf{y}}^{[m]}})$ be the Dirichlet form on $L^{2}(\mu_{\mathsf{a},R,\mathsf{y}}^{[m]})$ given in Lemma 10.2. Let $(\underline{\mathscr{E}}_{R\star}^{\mu_{\mathsf{a}}^{[m]}},\underline{\mathscr{D}}_{R\star}^{\mu_{\mathsf{a}}^{[m]}})$ be the superposition of $(\mathscr{E}_{R}^{\mu_{\mathsf{a},R,\mathsf{y}}^{[m]}},\underline{\mathscr{D}}_{R}^{\mu_{\mathsf{a},R,\mathsf{y}}^{[m]}})$ in $\mathsf{y}$ for $\mu_{\mathsf{a}}^{[m]}$, that is,
$\underline{\mathscr{E}}_{R\star}^{\mu_{\mathsf{a}}^{[m]}}(f,g)=\int_{\mathsf{S}}\mathscr{E}_{R}^{\mu_{\mathsf{a},R,\mathsf{y}}^{[m]}}(f,g)\mu_{\mathsf{a}}^{[m]}(d\mathsf{y}),$
$\underline{\mathscr{D}}_{R\star}^{\mu_{\mathsf{a}}^{[m]}}=\{f\in\bigcap_{\mathsf{y}\in\mathsf{S}_{0}}\underline{\mathscr{D}}_{R}^{\mu_{\mathsf{a},R,\mathsf{y}}^{[m]}};\underline{\mathscr{E}}_{R\star}^{\mu_{\mathsf{a}}^{[m]}}(f,f)<\infty,f\in L^{2}(\mu_{\mathsf{a}}^{[m]})\},$
where $\mathsf{S}_{0}$ is the set in Definition 2.3. We recall that $\mu(\mathsf{S}_{0})=1$. Applying the general theorem on the superposition of closed forms [4], we see that $(\underline{\mathscr{E}}_{R\star}^{\mu_{\mathsf{a}}^{[m]}},\underline{\mathscr{D}}_{R\star}^{\mu_{\mathsf{a}}^{[m]}})$ is a closed form on $L^{2}(\mu_{\mathsf{a}}^{[m]})$. Let $(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}}_{R},\underline{\mathscr{D}}_{R}^{\mu_{\mathsf{a}}^{[m]}})$ be as in Lemma 10.2.
Lemma 11.3.
Assume ((A1))((A3)). Let $m\in\{0\}\cup\mathbb{N}$. Then, for $\mu$-a.s. $\mathsf{a}$
$(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}}_{R},\underline{\mathscr{D}}_{R}^{\mu_{\mathsf{a}}^{[m]}})=(\underline{\mathscr{E}}_{R\star}^{\mu_{\mathsf{a}}^{[m]}},\underline{\mathscr{D}}_{R\star}^{\mu_{\mathsf{a}}^{[m]}})\quad\text{ for all }R\in\mathbb{N}.$
Proof.
The proof of Lemma 11.3 is identical to that of Lemma 11.1. Hence, it is omitted here. ∎
We easily show that $(\underline{\mathscr{E}}_{R\star}^{\mu_{\mathsf{a}}^{[m]}},\underline{\mathscr{D}}_{R\star}^{\mu_{\mathsf{a}}^{[m]}})$, $R\in\mathbb{N}$, is increasing. Let $(\underline{\mathscr{E}}_{\star}^{\mu_{\mathsf{a}}^{[m]}},\underline{\mathscr{D}}_{\star}^{\mu_{\mathsf{a}}^{[m]}})$ be the increasing limit of $(\underline{\mathscr{E}}_{R\star}^{\mu_{\mathsf{a}}^{[m]}},\underline{\mathscr{D}}_{R\star}^{\mu_{\mathsf{a}}^{[m]}})$ in the sense of (13.2). Then, $(\underline{\mathscr{E}}_{\star}^{\mu_{\mathsf{a}}^{[m]}},\underline{\mathscr{D}}_{\star}^{\mu_{\mathsf{a}}^{[m]}})$ is a closed form on $L^{2}(\mu_{\mathsf{a}}^{[m]})$ from Lemma 13.5 i.
Replacing $\mu$ by $\mu_{\mathsf{a}}$, we set $(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\underline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}}_{\bullet})$ similarly as $(\mathscr{E}^{\mu^{[m]}},\underline{\mathscr{D}}^{\mu^{[m]}}_{\bullet})$. Let $(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\underline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}})$ be the closed form on $L^{2}(\mu_{\mathsf{a}}^{[m]})$ in Lemma 10.6.
Lemma 11.4.
Assume ((A1))((A3)). Then, for each $m\in\{0\}\cup\mathbb{N}$,
$(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\underline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}}_{\bullet})$ $=(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\underline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}})=(\underline{\mathscr{E}}_{\star}^{\mu_{\mathsf{a}}^{[m]}},\underline{\mathscr{D}}_{\star}^{\mu_{\mathsf{a}}^{[m]}}).$ (11.5)
Proof.
Equation (11.5) follows from Lemma 10.5, Lemma 10.3 i, and Lemma 11.3. The detail of the proof is similar to that of Lemma 11.2. Hence, it is omitted here. ∎
Let $(\mathscr{E}^{\mu^{[m]}}_{\star},\overline{\mathscr{D}}^{\mu^{[m]}}_{\star})$, $(\underline{\mathscr{E}}_{\star}^{\mu^{[m]}},\underline{\mathscr{D}}_{\star}^{\mu^{[m]}})$, and $(\underline{\mathscr{E}}_{\star\star}^{\mu^{[m]}},\underline{\mathscr{D}}_{\star\star}^{\mu^{[m]}})$ be the superpositions of $(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\overline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}})$, $(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\underline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}})$, and $(\underline{\mathscr{E}}_{\star}^{\mu_{\mathsf{a}}^{[m]}},\underline{\mathscr{D}}_{\star}^{\mu_{\mathsf{a}}^{[m]}})$ in $\mathsf{a}$, respectively. Here, $(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\overline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}})$ is the closed form in Lemma 10.6. The following theorem asserts that these Dirichlet forms coincide and are strongly local and quasi-regular.
Theorem 11.1 (Uniqueness of extension of Dirichlet forms).
Assume ((A1))((A3)).
i Let $\mathsf{S}_{0}$ be the set in Definition 2.3. Then, for all $\mathsf{a}\in\mathsf{S}_{0}$,
$(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\overline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}})$ $=(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\underline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}})=(\underline{\mathscr{E}}_{\star}^{\mu_{\mathsf{a}}^{[m]}},\underline{\mathscr{D}}_{\star}^{\mu_{\mathsf{a}}^{[m]}}),$ (11.6)
$(\mathscr{E}^{\mu^{[m]}}_{\star},\overline{\mathscr{D}}^{\mu^{[m]}}_{\star})$ $=(\underline{\mathscr{E}}_{\star}^{\mu^{[m]}},\underline{\mathscr{D}}_{\star}^{\mu^{[m]}})=(\underline{\mathscr{E}}_{\star\star}^{\mu^{[m]}},\underline{\mathscr{D}}_{\star\star}^{\mu^{[m]}}).$ (11.7)
$i$ All Dirichlet forms in (11.6) and (11.7) are strongly local and quasi-regular.
Proof.
From Theorem 10.6 and Lemma 11.4, we obtain (11.6). Taking the superposition of (11.6) in $\mathsf{a}$, we have (11.7), which yields i.
From Lemma 10.6 ii, $(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\overline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}})$ is strongly local and quasi-regular on $L^{2}(\mu_{\mathsf{a}}^{[m]})$. Hence, all Dirichlet forms in (11.6) are strongly local and quasi-regular.
Let $\mathsf{a}]=\{\mathsf{b}\in\mathsf{S}_{0};\mu_{\mathsf{a}}=\mu_{\mathsf{b}}\}$ be as in Definition 2.3. Then, $\mu_{\mathsf{a}}([\mathsf{a}])=1$ and $\mathsf{a}]\cap[\mathsf{b}]=\emptyset$ if $\mathsf{a}]\neq[\mathsf{b}]$. Note that $\mathsf{a}]$ gives a partition of $\mathsf{S}_{0}$ and $\mu(\mathsf{S}_{0})=1$ by construction.
Let $\mathbf{S}_{\mathsf{a}}^{[m]}=(\mathbb{R}^{d})^{m}\times[\mathsf{a}]$. Then, $(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\overline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}})$ is a strongly local, quasi-regular Dirichlet form on $L^{2}(\mathbf{S}_{\mathsf{a}}^{[m]},\mu_{\mathsf{a}}^{[m]})$. Let $\mathrm{Cap}_{\mathsf{a}}^{[m]}$ be the capacity given by the Dirichlet form $(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\overline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}})$ on $L^{2}(\mathbf{S}_{\mathsf{a}}^{[m]},\mu_{\mathsf{a}})$. Then, there exists an increasing sequence of compact sets $\mathsf{K}_{\mathsf{a}}^{n}$ such that $\mathsf{K}_{\mathsf{a}}^{n}\subset[\mathsf{a}]$ satisfying
$\lim_{n\to\infty}\mathrm{Cap}_{\mathsf{a}}^{[m]}(\{(\mathbb{R}^{d})^{m}\times\mathsf{K}_{\mathsf{a}}^{n}\}^{c})=0.$ (11.8)
Indeed, (11.8) follows from ((QR1)) in Definition 13.4. Here, we take, for each $n\in\mathbb{N}$,
Let $\mathsf{P}^{\mu_{\mathsf{a}}^{[m]}}$ be the diffusion properly associated with the strongly local, quasi-regular Dirichlet form $(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\overline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}})$ on $L^{2}(\mu_{\mathsf{a}}^{[m]})$. From (11.8), the state space of $\mathsf{P}^{\mu_{\mathsf{a}}^{[m]}}$ can be taken to be disjoint for $\mathsf{a}]\neq[\mathsf{b}]$. Thus, the collection of diffusions $\mathsf{P}^{\mu_{\mathsf{a}}^{[m]}}$, $\mathsf{a}\in\mathsf{S}_{0}$, gives the diffusion associated with $(\mathscr{E}^{\mu^{[m]}}_{\star},\overline{\mathscr{D}}^{\mu^{[m]}}_{\star})$ on $L^{2}(\mu^{[m]})$. Hence, from [5, Th. 1.5.3], $(\mathscr{E}^{\mu^{[m]}}_{\star},\overline{\mathscr{D}}^{\mu^{[m]}}_{\star})$ on $L^{2}(\mu^{[m]})$ is quasi-regular. The strong locality of $(\mathscr{E}^{\mu^{[m]}}_{\star},\overline{\mathscr{D}}^{\mu^{[m]}}_{\star})$ is clear by construction. Hence, all Dirichlet forms in (11.7) are strongly local and quasi-regular on $L^{2}(\mu^{[m]})$. Thus, we complete the proof of ii. ∎
At first glance, Theorem 11.1 seems abstract, and it may not be easy to understand how it plays a role in the stochastic dynamics of an infinite-particle system. Therefore, to explain its importance, we describe some applications in the following remark.

### v4

@@ 11.1. Uniqueness of extensions of Dirichlet forms
Hence, the diffusion $\underline{\mathbf{X}}_{\mathsf{a},R,\mathsf{y}}^{[m]}$, $\mathsf{y}\in\mathsf{S}$, obtained from $(\mathscr{E}_{\mathsf{a},R}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R\star}^{[m]})$ and the diffusion $\underline{\mathbf{X}}_{\mathsf{a},R}^{[m]}$ associated with $(\mathscr{E}_{\mathsf{a},R,\mathsf{y}}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R,\mathsf{y}}^{[m]})$ coincide. This implies the identity of the Dirichlet forms in (11.2). ∎
@@
We see that $(\mathscr{E}_{\mathsf{a},R}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R\star}^{[m]})$, $R\in\mathbb{N}$, is increasing. Let $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{lwr}\star}^{[m]})$ be the increasing limit of $(\mathscr{E}_{\mathsf{a},R}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R\star}^{[m]})$. By Lemma 13.5 i, $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{lwr}\star}^{[m]})$ is a closed form on $L^{2}(\mu_{\mathsf{a}}^{[m]})$. Let $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{lwr}}^{[m]})$ be as in Lemma 10.4.
@@
For each $m\in\{0\}\cup\mathbb{N}$ and $\mu$-a.s. $\mathsf{a}$,
$(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{lwr}}^{[m]})$ $=(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{lwr}\star}^{[m]}).$ (11.3)
@@
Equations (11.3) follows from (11.2). ∎
Recall that $(\mathscr{E}_{\mathsf{a},R}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R\star}^{[m]})$ is a closed form on $L^{2}(\mu_{\mathsf{a},R,\mathsf{y}}^{[m]})$ and that $\mathscr{D}_{\circ\mathrm{b}}^{[m]}\subset\underline{\mathscr{D}}_{\mathsf{a},R,\mathsf{y}}^{[m]}$. Then $(\mathscr{E}_{\mathsf{a},R,\mathsf{y}}^{[m]},\mathscr{D}_{\circ\mathrm{b}}^{[m]})$ is closable on $L^{2}(\mu_{\mathsf{a},R,\mathsf{y}}^{[m]})$. Let $(\mathscr{E}_{\mathsf{a},R,\mathsf{y}}^{[m]},\overline{\mathscr{D}}_{\mathsf{a},R,\mathsf{y}}^{[m]})$ be its closure. Let $(\mathscr{E}_{\mathsf{a},R\star}^{[m]},\overline{\mathscr{D}}_{\mathsf{a},R\star}^{[m]})$ be the superposition of $(\mathscr{E}_{\mathsf{a},R,\mathsf{y}}^{[m]},\overline{\mathscr{D}}_{\mathsf{a},R,\mathsf{y}}^{[m]})$ with respect to $\mathsf{y}$ under $\mu$.
Let $(\mathscr{E}_{R}^{[m]},\overline{\mathscr{D}}_{\star,R}^{[m]})$ and $(\mathscr{E}_{R},\overline{\mathscr{D}}_{\star,R\star}^{[m]})$ be the superpositions of $(\mathscr{E}_{\mathsf{a},R}^{[m]},\overline{\mathscr{D}}_{\mathsf{a},R}^{[m]})$ and $(\mathscr{E}_{\mathsf{a},R}^{[m]},\overline{\mathscr{D}}_{\mathsf{a},R\star}^{[m]})$, respectively, with respect to $\mathsf{a}$ under $\mu$.
Lemma 11.3.
Let $m\in\{0\}\cup\mathbb{N}$ and $R\in\mathbb{N}$. Then the following hold:
$(\mathscr{E}_{\mathsf{a},R}^{[m]},\overline{\mathscr{D}}_{\mathsf{a},R}^{[m]})=(\mathscr{E}_{\mathsf{a},R}^{[m]},\overline{\mathscr{D}}_{\mathsf{a},R\star}^{[m]}),$ (11.4)
$(\mathscr{E}_{R}^{[m]},\overline{\mathscr{D}}_{\star,R}^{[m]})=(\mathscr{E}_{R}^{[m]},\overline{\mathscr{D}}_{\star,R\star}^{[m]}).$ (11.5)
Proof.
Since the carré du champ of $\mathscr{E}_{R}^{[m]}$ vanishes outside $\overline{S}_{R}^{c}$, we obtain (11.4). Equation (11.5) follows immediately from (11.4). ∎
Let $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}}^{[m]})$ be the closed form given in Lemma 10.4. Let $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}\star}^{[m]})$ be the decreasing limit of $(\mathscr{E}_{\mathsf{a},R}^{[m]},\overline{\mathscr{D}}_{\mathsf{a},R\star}^{[m]})$.
Lemma 11.4.
Let $m\in\{0\}\cup\mathbb{N}$ and $R\in\mathbb{N}$. Then the following hold:
$(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}}^{[m]})=(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}\star}^{[m]}).$ (11.6)
Proof.
(11.6) follows immediately from Lemma 10.4 and (11.5). ∎
Let $(\mathscr{E}^{[m]},\mathscr{D}_{\star,\mathrm{upr}}^{[m]})$, $(\mathscr{E}^{[m]},\mathscr{D}_{\star,\mathrm{upr}\star}^{[m]})$, $(\mathscr{E}^{[m]},\mathscr{D}_{\star,\mathrm{lwr}}^{[m]})$, and $(\mathscr{E}^{[m]},\mathscr{D}_{\star,\mathrm{lwr}\star}^{[m]})$ be the superpositions of $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}}^{[m]})$, $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}\star}^{[m]})$, $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{lwr}}^{[m]})$, and $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{lwr}\star}^{[m]})$, respectively, with respect to $\mathsf{a}$ under $\mu$.
$\mathscr{E}^{[m]}(f,g)$ $=\int_{\mathsf{S}}\mathscr{E}_{\mathsf{a}}^{[m]}(f,g)\,\mu(d\mathsf{a}),$
$\mathscr{D}_{\star,\mathrm{upr}}^{[m]}$ $=\Bigl\{f\in\bigcap_{\mu\text{-a.s.\,}\mathsf{a}}\mathscr{D}_{\mathsf{a},\mathrm{upr}}^{[m]}\,;\,\mathscr{E}^{[m]}(f,f)<\infty,\,f\in L^{2}(\mu^{[m]})\Bigr\}.$
Replacing $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}}^{[m]})$ with $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}*}^{[m]})$, $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{lwr}}^{[m]})$, and $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{lwr}\star}^{[m]})$, we obtain the superpositions $(\mathscr{E}^{[m]},\mathscr{D}_{\star,\mathrm{upr}\star}^{[m]})$, $(\mathscr{E}^{[m]},\mathscr{D}_{\star,\mathrm{lwr}}^{[m]})$ and $(\mathscr{E}^{[m]},\mathscr{D}_{\star,\mathrm{lwr}\star}^{[m]})$, respectively.
The following theorem shows that these Dirichlet forms coincide and are strongly local and quasi-regular.
Theorem 11.1 (Uniqueness of extensions of Dirichlet forms).
i Let $\mathsf{S}_{0}$ be the set in Definition 2.2. For $\mu$-a.s. $\mathsf{a}\in\mathsf{S}_{0}$,
$(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}}^{[m]})=(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}\star}^{[m]})=(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{lwr}}^{[m]})=(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{lwr}\star}^{[m]}),$ (11.7)
$(\mathscr{E}^{[m]},\mathscr{D}_{\star,\mathrm{upr}}^{[m]})=(\mathscr{E}^{[m]},\mathscr{D}_{\star,\mathrm{upr}\star}^{[m]})=(\mathscr{E}^{[m]},\mathscr{D}_{\star,\mathrm{lwr}\star}^{[m]})=(\mathscr{E}^{[m]},\mathscr{D}_{\star,\mathrm{lwr}\star}^{[m]}).$ (11.8)
$i$ All Dirichlet forms in (11.7) and (11.8) are strongly local and quasi-regular on $L^{2}(\mu_{\mathsf{a}}^{[m]})$ and $L^{2}(\mu^{[m]})$, respectively.
Proof.
From (11.6), (10.9) and (11.3), we obtain (11.7). Taking the superposition of (11.7) with respect to $\mathsf{a}$ under $\mu$ yields (11.8). This proves i.
By Lemma 10.7 ii, $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}}^{[m]})$ is strongly local and quasi-regular on $L^{2}(\mu_{\mathsf{a}}^{[m]})$. Hence, all Dirichlet forms in (11.7) are strongly local and quasi-regular.
Let $\mathsf{a}]=\{\mathsf{b}\in\mathsf{S}_{0};\mu_{\mathsf{a}}=\mu_{\mathsf{b}}\}$ be as in Definition 2.2. We have $\mu_{\mathsf{a}}([\mathsf{a}])=1$, and $\mathsf{a}]\cap[\mathsf{b}]=\emptyset$ if $\mathsf{a}]\neq[\mathsf{b}]$. Note that $\mathsf{a}]$ gives a partition of $\mathsf{S}_{0}$ and $\mu(\mathsf{S}_{0})=1$ by construction.
Let $\mathrm{Cap}_{\mathsf{a}}^{[m]}$ be the capacity given by $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}}^{[m]})$ on $L^{2}((\mathbb{R}^{d})^{m}\times[\mathsf{a}],\mu_{\mathsf{a}})$. There exists an increasing sequence of compact sets $\mathsf{K}_{\mathsf{a}}^{n}$ satisfying
$\mathsf{K}_{\mathsf{a}}^{n}\subset[\mathsf{a}],\quad\lim_{n\to\infty}\mathrm{Cap}_{\mathsf{a}}^{[m]}(\{(\mathbb{R}^{d})^{m}\times\mathsf{K}_{\mathsf{a}}^{n}\}^{c})=0.$ (11.9)
Indeed, (11.9) follows from ((Q1)) in Definition 13.4. Here, we take, for each $n\in\mathbb{N}$,
Let $\mathsf{P}_{\mathsf{a}}^{[m]}$ be the diffusion properly associated with the Dirichlet form $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}}^{[m]})$ on $L^{2}(\mu_{\mathsf{a}}^{[m]})$, which is strongly local and quasi-regular by Lemma 10.7 ii. From (11.9), the state space of $\mathsf{P}_{\mathsf{a}}^{[m]}$ can be taken to be disjoint for $\mathsf{a}]\neq[\mathsf{b}]$. Thus, the collection of diffusions $\mathsf{P}_{\mathsf{a}}^{[m]}$, $\mathsf{a}\in\mathsf{S}_{0}$, gives the diffusion properly associated with $(\mathscr{E}^{[m]},\mathscr{D}_{\star,\mathrm{upr}}^{[m]})$ on $L^{2}(\mu^{[m]})$.
Hence, from [5, Th. 1.5.3], $(\mathscr{E}^{[m]},\mathscr{D}_{\star,\mathrm{upr}}^{[m]})$ on $L^{2}(\mu^{[m]})$ is quasi-regular. The strong locality of $(\mathscr{E}^{[m]},\mathscr{D}_{\star,\mathrm{upr}}^{[m]})$ is clear by construction. Thus, all Dirichlet forms in (11.8) are strongly local and quasi-regular on $L^{2}(\mu^{[m]})$. This completes the proof of ii. ∎
Theorem 11.1 plays a crucial role in the proof of Theorems 2.42.5. We present further applications of Theorem 11.1 in the following remark.
@@
i The sub-diffusivity of tagged particles in the Ginibre interacting Brownian motion ($d=\beta=2$) follows from (11.7) [31], reflecting the long-range nature of the two-dimensional Coulomb potential. In contrast, tagged particles are always diffusive for Ruelle-class potentials with a convex hardcore in dimensions $d\geq 2$ [25].
ii Suzuki proved that the unlabeled Markov processes associated with a Dirichlet form are ergodic if the reference measure $\mu$ is number rigid and the Dirichlet form is locally ergodic [40]. This theorem can be applied to the Ginibre RPF. In [40], Suzuki considered the Dirichlet form that corresponds to $(\mathscr{E}^{[m]},\mathscr{D}_{\star,\mathrm{lwr}\star}^{[m]})$. By Theorem 11.1, the continuous Markov process considered in [40] is an ergodic diffusion.
iii Suzuki [41] studied curvature bounds for Dyson’s Brownian motion, an infinite-particle system on $\mathbb{R}$ with the two-dimensional Coulomb interaction at inverse temperature $\beta=2$, that is, with a Riesz potential. He analyzed Dirichlet forms of the type $(\mathscr{E}^{[m]},\mathscr{D}_{\star,\mathrm{lwr}\star}^{[m]})$. Since our argument for Theorem 11.1 may be valid for Riesz potentials, his results can apply to Dyson’s model studied in [27, 34] by the uniqueness of Dirichlet forms.